Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
capocasa, ruggi - 23:32 Thursday 05 May 2022 (55726) Print this report
Comment to PR centering (55725)

During the beam shift activity, we took the chance to check the newly implemented channels to monitor beam shift on NI/WI.

Dithering lines generated in the Acl_pre process are sent to input marionettes with the following frequencies: WI/NI_TX: 1.5 Hz, WI/NI_TY: 1.3 Hz.

Then B7/B8 are demodulated at the frequency of the lines providing the monitoring signals ASC_{NI,WI}_{TX,TY}_LF_{B7,B8}_{I,Q}_FS

Fig.1 shows such demodulated signals (I and Q), accounting for NI/WI decentering (both horizonal and vertical), during the beam shift.

The demodulation phase has been tuned offline by taking atan(I/Q).

Fig 2. shows the magnitude of the signals. We plan to do a proper calibration of the signals and turn them on on regular basis to monitor the beam shift.

A function to switch on/off the lines has been implemented on

{WI,NI}_LF_line_toggle (ampl_TX,ampl_TY). The amplitude used today was 1e2.

Images attached to this comment
Search Help